# Circular strings, magnons, plane waves and local quenches in BTZ

### v1

@@ Abstract
We show that string theory on the geometry $BTZ\times S^{3}\times M$ supported with either Neveu-Schwarz flux or Ramond flux admits states which obey identical dispersion relations to those of classical solutions like circular strings, giant magnons, or plane wave excitations in the geometry $AdS_{3}\times S^{3}\times M$. Here, $M$ can be $T^{4}$, $K3$, or $S^{3}\times S^{1}$. This is made possible by the map, which takes the particle at the origin of $AdS_{3}$ with angular momentum along one of the angles of $S_{3}$ to a particle falling into the BTZ horizon. We use this map to construct circular strings, magnons, as well as plane waves in the BTZ geometry. We show that the $SL(2,R)$ charges of these states on $AdS_{3}$ and that of the corresponding states in the BTZ geometry are related by a boost. The dual description of these states in the BTZ geometry are local quench in the thermal CFT. These quenches carry energy density, $R$-charges, non-trivial expectation value of the marginal operator dual to the dilaton and move on the light cone in CFT. In general, the left and the right moving quenches are not symmetric.
@@ 1 Introduction
The BTZ black hole Banados:1992wn ; Banados:1992gq It is one of the most well-studied solutions among black hole geometries. It also occurs as the near-horizon geometry of the $D1D5$ black hole in string theory Maldacena:1998bw . It serves as a prototype for studies of black hole entropy, information paradox, and the holographic dual of large $c$ CFTs at finite temperature. This is partly because many aspects of the BTZ black hole are exactly solvable. Solutions for geodesics, wave equations of arbitrary spin fields Datta:2011za ; Datta:2012gc can be obtained in closed form as it is locally $AdS_{3}$. String propagation has also been studied earlier using the WZW model of the BTZ geometry. Natsuume:1996ij ; Satoh:1997xe ; Maldacena:2000hw ; Maldacena:2000kv ; Hemming:2001we ; Hemming:2002kd . However, a clear physical interpretation of the states of the string is lacking. This is important due to the existence of the horizon. Clearly, there must be string states that fall into the horizon, and these have not been identified and studied in earlier work.
@@
Finally, we examine the interpretation of the states that we have found in the BTZ geometry. These states should be excitations of the thermal state of the corresponding boundary CFT. The boundary CFT in this case is the symmetric product CFT on $T^{4}$ or $K3$, or in the case $S^{3}\times S^{1}$, it is the instanton moduli space on $S^{3}\times S^{1}$. We show that the classical states in the bulk, geodesics, circular strings, and magnons are dual to local quenches similar to that studied in Caputa:2014eta for the 3 dimensional BTZ geometry which carry only energy density and move on the light cone in the CFT. Here, however, the quenches in addition to energy density carry $R$ charges which arise due to the angular momentum on $S^{3}$. These states also have a non-trivial expectation value of the marginal operator dual to the dilaton in the CFT. We show that for the more general case of the map in (1), which allows transverse velocity of the in-falling particle in the $x$ direction, the dual quenches on the left and right light cones are not symmetric, and the amplitude of the pulse differs. Thus, these quenches generalise those found first in Caputa:2014eta . The figure 1 shows the plot of the energy density of these quenches, which clearly shows the asymmetric left and right moving pulses travelling at the speed of light.
@@ 2.1 The in-falling particle in BTZ
where the $\Sigma$’s are relate to the Pauli matrices $\hat{\sigma}_{i}$ as
@@
Let us evaluate the $SL(2,R)$ charges on the time like a geodesic sitting at the origin described by the trajectory (8). Note that for geodesics, the contribution proportional to $q$ in the charges (25) vanishes. The left and right $SL(2,R)$ charges for this geodesic are given by
@@
$Q_{L}\cdot Q_{L}|_{\tiny AdS_{3}}=-\kappa^{2}$ (67)
@@ 3 In-falling classical strings in BTZ
In $AdS_{3}\times S^{3}\times M$ several classical string solutions are constructed as follows: One considers the time like a geodesic at the origin of $AdS_{3}$, and a classical solution of the sigma model on $S^{3}$ such that the solution satisfies the classical Virasoro condition. To proceed further, we describe some aspects of the $SU(2)$ WZW model, it is given by the action
@@
$J_{L}^{i}=ik\ {\rm Tr}[\mathcal{\varsigma}^{i}.\hat{Q}_{L}],\quad J_{R}^{i}=ik\ {\rm Tr}[\mathcal{\varsigma}^{i}.\hat{Q}_{R}],$ (109)
@@
The 2nd Virasoro constraint in (117) is satisfied, since we still do not have world sheet spatial dependence in the $BTZ$ and the world sheet momentum of each of these solutions vanish on $S^{3}$ and $M$. Finally, the first Virasoro constraint is satisfied due to the non-trivial observation in the previous section. The maps (2.1), (2.2), which take the time-like geodesic at the origin of $AdS_{3}$ to the in-falling geodesic in BTZ is such that the $SL(2,R)$ Casimir $\kappa^{2}$ is invariant as shown for example in (2.1), (2.1). The world sheet stress tensor in BTZ on solutions which depend only on world sheet time is given by
@@
For completeness, we mention that the boundary conditions the solution in (128) satisfies in the limit (131) is given by
@@
We can also generalise the embedding for the giant magnons solutions in the presence of NS flux $q\neq 0$, found in Hoare:2013lja . All we have to do is take the BTZ part of the solution to be the in-falling geodesic obeying the trajectory (79), (81), (89) and the $S^{3}$ part to be that constructed in the earlier works.
@@ 4 Plane wave spectra in BTZ
[…]
@@ 4.1 The Penrose limit of $BTZ\times S^{3}\times S^{3}\times S^{1}$
In this section, we obtain the plane wave limit $BTZ\times S^{3}\times S^{3}\times S^{1}$ on in-falling geodesics. We choose to study the case of $M=S^{3}\times S^{1}$ with mixed Ramond-Ramond and NS fluxes with angular momentum on both $S^{3}$’s. Studying this general case enables the possibility of taking limits so that we can obtain a situation with either pure Ramond-Ramond flux or only pure NS flux, and with angular momentum only on one of the $S^{3}$. We will compare our result for the plane wave geometry with the plane wave limit of $AdS_{3}\times S^{3}\times S^{3}\times S^{1}$ with mixed fluxes obtained in Dei:2018yth . The gravity background of $BTZ\times S^{3}\times S^{3}\times S^{2}$ is given by
@@
The reason for the normalisations in the angular velocity is that the angular momentum from the Noether definition is defined as $J_{1}$. This can be seen as follows, from the Noether definition of the charge corresponding to shifts in $\psi_{1}$, WZW model on $S^{2}$ at level $k_{1}$, the angular momentum is given by
@@ 5 In-falling strings as local quenches
To obtain the last equality, we have used the Brown-Henneaux formula
@@
$\rho_{\epsilon}={\cal N}e^{-itHt}\Big(e^{-\epsilon H}{\cal O}(0,0)e^{\epsilon H}\Big)\rho_{\beta}\Big(e^{-\epsilon H}{\cal O}^{\dagger}(0,0)e^{\epsilon H}\Big)e^{iHt},$
@@ 5.1 CFT duals of geodesics solutions in $AdS_{3}\times S^{3}\times M$
Let us dimensionally reduce these sources on $AdS_{3}\times S^{3}$ and focus on the lowest Kaluza-Klein mode on the sphere $S^{3}$. This can be done by integrating out the coordinates on $S^{3}\times M$. Therefore, dimensional reduction of the stress tensor components in (199) on $S^{3}\times M^{3}$, we obtain the following sources
@@
$\frac{1}{\sqrt{g}}\epsilon^{\mu\nu\rho}\partial_{\nu}A_{\rho}=16\pi G_{N}\hat{J}^{\nu},\qquad$
@@
$\frac{\langle{\cal O}|T_{\phi\phi}(0,0)|{\cal O}\rangle}{\langle{\cal O}|{\cal O}\rangle}=\frac{\langle{\cal O}|T_{zz}(0)-T_{\bar{z}\bar{z}}(0)|{\cal O}\rangle}{\langle{\cal O}|{\cal O}\rangle}$ $=$ $0.$ (225)
@@ 5.2 In falling spinning geodesics in $BTZ\times S^{3}\times M$ as quenches
Observe that the background $AdS_{3}$ in (232) gets converted to BTZ, while the correction proportional to $\tilde{m}$ will be written as the back reaction to the BTZ background. The expressions for the perturbation are long and cumbersome, but as we have kept the leading order at the boundary, whose boundary coordinates are $\tau,x$. We can now go over to Fefferman-Graham coordinates and read out the expectation value of the boundary stress tensor. This is given by
@@
To get an idea of the width of the pulses, it is convenient to take the limit $\beta>>x,\tau$ and $\eta_{1}>>1,\eta\sim 1$, then the widths of the two pulses are given by
@@
$\langle T_{\tau\tau}\rangle_{CFT}$ $=$ ${\rm Tr}\big[\rho T_{tt}(x-\tau,x+\tau)\big],$
$=$ $\frac{\langle{\cal O}^{\dagger}(w_{2},\bar{w}_{2})T_{tt}(x-\tau,x+\tau){\cal O}(w_{1},\bar{w}_{1})\rangle_{\beta}}{\langle{\cal O}^{\dagger}(w_{4},\bar{w}_{4}){\cal O}(w_{1},\bar{w}_{1})\rangle_{\beta}}.$
@@ 5.3 Quenches dual to the circular string
$T_{\mu\nu}=\frac{\Delta}{R}\frac{\delta(r)\delta(\phi)}{r}.$ (270)
@@ Acknowledgements.
We thank Rajesh Gopakumar for discussions at various stages of this project, whose questions led us to study the spacetime charges of the in-falling states in BTZ. Rahul Metya would like to thank Ritwick Kumar Ghosh for stimulating discussions. J.R.D thanks S. Prem Kumar and Suvajit Majumder for an enjoyable collaboration during the difficult Covid times in which asymmetric quenches were first found in the BTZ background.
@@ Appendix A Penrose limit of Schwarzschild black hole in $AdS_{5}\times S^{5}$
$R^{-2}ds^{2}=-\left(r^{2}+1-\frac{r_{0}^{2}}{r^{2}}\right)dt^{2}+\frac{dr^{2}}{\left(r^{2}+1-\frac{r_{0}^{2}}{r^{2}}\right)}+r^{2}\left(d\phi^{2}+\sin^{2}{\phi}\ d\Omega^{2}\right)+d\psi^{2}+sin^{2}\psi\ d\Omega^{2}_{4}$ (273)
@@
$g_{\lambda\lambda}=g_{\lambda\theta}=0,\quad\text{and}\quad g_{\lambda\xi}=1$ (274)
@@
$t(\lambda,\xi,\theta)=\int_{0}^{\lambda}\dot{t}\ d\lambda^{\prime}-\frac{\xi}{E}+J\theta\ ,\quad r(\lambda,\xi,\theta)=\int_{0}^{\lambda}\dot{r}\ d\lambda^{\prime}\ ,\quad\phi(\lambda,\xi,\theta)=E\theta+\int_{0}^{\lambda}\dot{\phi}d\lambda^{\prime}$ (275)
@@
$+\left(-a_{0}J\dot{t}+JE\right)2d\lambda d\theta+(a_{0}JE^{-1})2d\xi d\theta+r^{2}\sin^{2}{\phi}\ d\Omega^{2}_{2}+ds^{2}(S^{5})$
$=2d\lambda d\xi-a_{0}E^{-2}d\xi^{2}+\left(r^{2}E^{2}-J^{2}a_{0}\right)d\theta^{2}+Ja_{0}E^{-1}2d\xi d\theta+r^{2}\sin^{2}{\phi}\ d\Omega^{2}_{2}+ds^{2}(S^{5})$
@@
$\lambda\rightarrow\lambda,\quad\xi\rightarrow\frac{\xi}{R^{2}},\quad\theta\rightarrow\frac{\theta}{R},\quad\psi\rightarrow\frac{\psi}{R}$ (276)
@@
$ds^{2}$ $=2d\lambda d\xi+\left(r^{2}E^{2}-J^{2}a_{0}\right)d\theta^{2}+r^{2}\sin^{2}\left(\int_{0}^{\lambda}\frac{J}{r^{2}}d\lambda^{\prime}\right)\ ds^{2}(\mathbb{R}^{2})+ds^{2}(\mathbb{R}^{5})$ (277)
$=2d\lambda d\xi+{\sum\limits_{i=1}^{8}\left(C_{i}dy^{i}\right)^{2}}$ (278)
@@
$\lambda=x^{+},\quad\xi=x^{-}+\frac{1}{2}\sum\limits_{i}\frac{1}{C_{i}}\left(\frac{\partial C_{i}}{\partial x^{+}}\right)x^{i}x^{i},\quad y^{i}=C^{-1}_{i}x^{i}$ (279)
@@
$=2dx^{+}dx^{-}+\sum\limits_{i}\left(M_{ii}x^{i}x^{i}\right)\left(dx^{+}\right)^{2}+\sum\limits_{i}dx^{i}dx^{i}$
@@
$\frac{d^{2}}{d\lambda^{2}}=\frac{d}{d\lambda}\left(\dot{r}\frac{d}{dr}\right)=\dot{r}^{2}\frac{d^{2}}{dr^{2}}+\dot{r}\left(\frac{d\dot{r}}{dr}\right)\frac{d}{dr}$ (280)
@@
$M_{11}$ $=r^{-1}\frac{d}{dr}\left[\dot{r}\frac{d}{dr}\left(\dot{r}r\right)\right]=\frac{4J^{2}r_{0}^{2}}{r^{6}};\quad M_{44}=...=M_{88}=0$
$M_{22}$ $=M_{33}=\frac{\dot{r}}{r\ \sin{\left(\int\frac{J}{r^{2}}du\right)}}\frac{d}{dr}\left[\dot{r}\frac{d}{dr}\left\{r\sin{\left(\int\frac{J}{r^{2}}du\right)}\right\}\right]=-\frac{2J^{2}r_{0}^{2}}{r^{6}}$ (281)
@@
$ds^{2}=2dx^{+}dx^{-}+\left[\frac{4J^{2}r_{0}^{2}}{r^{6}}x_{1}^{2}-\frac{2J^{2}r_{0}^{2}}{r^{6}}(x_{2}^{2}+x_{3}^{2})\right]\left(dx^{+}\right)^{2}+\sum\limits_{i=1}^{3}dx_{i}^{2}+\sum\limits_{i=1}^{5}dy_{i}^{2}$ (282)
@@
Now expanding the metric in the large R limit, we obtain the pp wave in the Rosen Coordinate

### v2

@@ Abstract
We show that string theory on the geometry $BTZ\times S^{3}\times M$ supported with either Neveu-Schwarz flux or Ramond flux admits states which obey identical dispersion relations to those of classical solutions like circular strings, giant magnons, or plane wave excitations in the geometry $AdS_{3}\times S^{3}\times M$. Here, $M$ can be $T^{4}$, $K3$, or $S^{3}\times S^{1}$. This is made possible by the map, which takes the particle at the origin of $AdS_{3}$ with angular momentum along one of the angles of $S_{3}$ to a particle falling into the BTZ horizon. We use this map to construct circular strings, magnons, as well as plane waves in the BTZ geometry. We show that the $SL(2,R)$ charges of these states on $AdS_{3}$ and that of the corresponding states in the BTZ geometry are related by a boost. The dual description of these states in the BTZ geometry are local quenches in the thermal CFT. These quenches carry energy density, $R$-charges, non-trivial expectation value of the marginal operator dual to the dilaton and move on the light cone in CFT. In general, the left and the right moving quenches are not symmetric.
@@ 1 Introduction
The BTZ black hole Banados:1992wn ; Banados:1992gq , It is one of the most well-studied solutions among black hole geometries. It also occurs as the near-horizon geometry of the $D1D5$ black hole in string theory Maldacena:1998bw . It serves as a prototype for studies of black hole entropy, information paradox, and the holographic dual of large $c$ CFTs at finite temperature. This is partly because many aspects of the BTZ black hole are exactly solvable. Solutions for geodesics, wave equations of arbitrary spin fields Datta:2011za ; Datta:2012gc can be obtained in closed form as it is locally $AdS_{3}$. String propagation has also been studied earlier using the WZW model of the BTZ geometry. Natsuume:1996ij ; Satoh:1997xe ; Maldacena:2000hw ; Maldacena:2000kv ; Hemming:2001we ; Hemming:2002kd . However, a clear physical interpretation of the states of the string is lacking. This is important due to the existence of the horizon. Clearly, there must be string states that fall into the horizon, and these have not been identified and studied in earlier work.
@@
Finally, we examine the interpretation of the states that we have found in the BTZ geometry. These states should be excitations of the thermal state of the corresponding boundary CFT. The boundary CFT in this case is the symmetric product CFT on $T^{4}$ or $K3$, or in the case $S^{3}\times S^{1}$, it is the instanton moduli space on $S^{3}\times S^{1}$. We show that the classical states in the bulk, geodesics, circular strings, and magnons are dual to local quenches similar to that studied in Caputa:2014eta for the 3 dimensional BTZ geometry which carry only energy density and move on the light cone in the CFT. Here, however, the quenches in addition to energy density carry $R$ charges which arise due to the angular momentum on $S^{3}$. These states also have a non-trivial expectation value of the marginal operator dual to the dilaton in the CFT. We show that for the more general case of the map in (1), which allows transverse velocity of the in-falling particle in the $x$ direction, the dual quenches on the left and right light cones are not symmetric, and the amplitude of the pulse differs. Thus, these quenches generalise those found first in Caputa:2014eta . The figure 1 shows the plot of the energy density of these quenches, which clearly shows the asymmetric left and right-moving pulses travelling at the speed of light.
@@ 2.1 The in-falling particle in BTZ
where the $\Sigma$’s are related to the Pauli matrices $\hat{\sigma}_{i}$ as
@@
Let us evaluate the $SL(2,R)$ charges on the time like geodesic sitting at the origin described by the trajectory (8). Note that for geodesics, the contribution proportional to $q$ in the charges (25) vanishes. The left and right $SL(2,R)$ charges for this geodesic are given by
@@
$Q_{L}\cdot Q_{L}|_{\tiny AdS_{3}}=-\kappa^{2}.$ (67)
@@ 3 In-falling classical strings in BTZ
In $AdS_{3}\times S^{3}\times M$ several classical string solutions are constructed as follows: One considers a time like geodesic at the origin of $AdS_{3}$, and a classical solution of the sigma model on $S^{3}$ such that the solution satisfies the classical Virasoro condition. To proceed further, we describe some aspects of the $SU(2)$ WZW model, it is given by the action
@@
$J_{L}^{i}=i\ {\rm Tr}[\mathcal{\varsigma}^{i}.\hat{Q}_{L}],\quad J_{R}^{i}=i\ {\rm Tr}[\mathcal{\varsigma}^{i}.\hat{Q}_{R}],$ (109)
@@
The 2nd Virasoro constraint in (117) is satisfied, since we still do not have world sheet spatial dependence in the $BTZ$ and the world sheet momentum of each of these solutions vanish on $S^{3}$ and $M$. Finally, the first Virasoro constraint is satisfied due to the non-trivial observation in the previous section. The maps (2.1), (2.2), which take the time-like geodesic at the origin of $AdS_{3}$ to the in-falling geodesic in BTZ is such that the $SL(2,R)$ Casimir $\kappa^{2}$ is invariant as shown for example in (2.1), (2.1). The world sheet stress tensor in BTZ on solutions, which depends only on world sheet time, is given by
@@
For completeness, we mention that the boundary conditions that the solution in (128) satisfies in the limit (131) is given by
@@
We can also generalise the embedding for the giant magnon solutions in the presence of NS flux $q\neq 0$, found in Hoare:2013lja . All we have to do is take the BTZ part of the solution to be the in-falling geodesic obeying the trajectory (79), (81), (89) and the $S^{3}$ part to be that constructed in the earlier works.
@@ 4 Plane wave spectra in BTZ
In section 3, we demonstrated that the classical solutions of the string sigma model $AdS_{3}\times S^{3}\times M$ can be mapped to solutions in $BTZ\times S^{3}\times M$. These solutions involve the timelike geodesics at the origin of $AdS_{3}$, which are mapped to an in-falling geodesic in BTZ. We also showed that the solutions obey an identical dispersion relation once the conformal dimension $\Delta^{2}$ is identified with the quadratic Casimir of the WZW model on BTZ. The plane wave limit of $AdS_{3}\times S^{3}\times M$ Berenstein:2002jq , involves a time-like geodesic at the origin with angular momentum along one of the circles of $S^{3}$. The observations in section 3 suggest that it should be possible to obtain the same plane wave geometry in $BTZ\times S^{3}\times M$ by zooming in on the geometry around in-falling geodesics in BTZ. Contrary to the above argument, it is known that Penrose limit around generic geodesics in black hole backgrounds leads to time-dependent singular plane waves Blau:2003dz ; Blau:2004yi . Let us briefly review these results to place the observation for the Penrose limit for in-falling geodesics in the BTZ black hole in a relevant context. Consider the Schwarzschild black hole in $AdS_{5}\times S^{5}$, the metric is given in
@@ 4.1 The Penrose limit of $BTZ\times S^{3}\times S^{3}\times S^{1}$
In this section, we obtain the plane wave limit $BTZ\times S^{3}\times S^{3}\times S^{1}$ on in-falling geodesics. We choose to study the case of $M=S^{3}\times S^{1}$ with mixed Ramond-Ramond and NS fluxes with angular momentum on both $S^{3}$’s. Studying this general case enables the possibility of taking limits so that we can obtain a situation with either pure Ramond-Ramond flux or only pure NS flux, and with angular momentum only on one of the $S^{3}$. We will compare our result for the plane wave geometry with the plane wave limit of $AdS_{3}\times S^{3}\times S^{3}\times S^{1}$ with mixed fluxes obtained in Dei:2018yth . The gravity background of $BTZ\times S^{3}\times S^{3}\times S^{1}$ is given by
@@
The reason for the normalisations in the angular velocity is that the angular momentum from the Noether definition is defined as $J_{1}$. This can be seen as follows, from the Noether definition of the charge corresponding to shifts in $\psi_{1}$, WZW model on $S^{3}$ at level $k_{1}$, the angular momentum is given by
@@ 5 In-falling strings as local quenches
To obtain the last equality, we have used the Brown-Henneaux formula Brown:1986nw
@@
$\rho_{\epsilon}={\cal N}e^{-iHt}\Big(e^{-\epsilon H}{\cal O}(0,0)e^{\epsilon H}\Big)\rho_{\beta}\Big(e^{-\epsilon H}{\cal O}^{\dagger}(0,0)e^{\epsilon H}\Big)e^{iHt},$
@@ 5.1 CFT duals of geodesics solutions in $AdS_{3}\times S^{3}\times M$
Let us dimensionally reduce these sources on $AdS_{3}\times S^{3}$ and focus on the lowest Kaluza-Klein mode on the sphere $S^{3}$. This can be done by integrating out the coordinates on $S^{3}\times M$. Therefore, dimensional reduction of the stress tensor components in (199) on $S^{3}\times M$, we obtain the following sources
@@
$\frac{R}{\sqrt{g}}\epsilon^{\mu\nu\rho}\partial_{\nu}A_{\rho}=16\pi G_{N}\hat{J}^{\nu},\qquad$
@@
$\frac{\langle{\cal O}|T_{\phi\phi}(0,0)|{\cal O}\rangle}{\langle{\cal O}|{\cal O}\rangle}=\frac{\langle{\cal O}|T_{ww}(0)+T_{\bar{w}\bar{w}}(0)|{\cal O}\rangle}{\langle{\cal O}|{\cal O}\rangle}$ $=$ $\frac{\Delta}{R^{2}}-\frac{c}{12R^{2}}.$ (225)
@@ 5.2 In falling spinning geodesics in $BTZ\times S^{3}\times M$ as quenches
Observe that the background $AdS_{3}$ in (232) gets converted to BTZ, while the correction proportional to $\tilde{m}$ will be written as the back reaction to the BTZ background. The expressions for the perturbation are long and cumbersome, but we have kept the leading order at the boundary, whose boundary coordinates are $\tau,x$. We can now go over to Fefferman-Graham coordinates and read out the expectation value of the boundary stress tensor. This is given by
@@
To get an idea of the width of the pulses, it is convenient to take the limit $\beta>>x,\tau$ and $\eta_{1}>>1,\eta_{2}\sim 1$, then the widths of the two pulses are given by
@@
$\langle T_{\tau\tau}\rangle_{CFT}$ $=$ ${\rm Tr}\big[\rho T_{\tau\tau}(x-\tau,x+\tau)\big],$
$=$ $\frac{\langle{\cal O}^{\dagger}(w_{2},\bar{w}_{2})T_{\tau\tau}(x-\tau,x+\tau){\cal O}(w_{1},\bar{w}_{1})\rangle_{\beta}}{\langle{\cal O}^{\dagger}(w_{4},\bar{w}_{4}){\cal O}(w_{1},\bar{w}_{1})\rangle_{\beta}}.$
@@ 5.3 Quenches dual to the circular string
$T^{tt}=\frac{\Delta}{R}\frac{\delta(r)\delta(\phi)}{r}.$ (270)
@@ Acknowledgements.
We thank Rajesh Gopakumar for discussions at various stages of this project, whose questions led us to study the spacetime charges of the in-falling states in BTZ. RM would like to thank Ritwick Kumar Ghosh for stimulating discussions. J.R.D thanks S. Prem Kumar and Suvajit Majumder for an enjoyable collaboration during the difficult Covid times in which asymmetric quenches were first found in the BTZ background. RM acknowledge support of the Department of Atomic Energy, Government of India, under project no. RTI4019.
@@ Appendix A Penrose limit of Schwarzschild black hole in $AdS_{5}\times S^{5}$
$R^{-2}ds^{2}=-\left(r^{2}+1-\frac{r_{0}^{2}}{r^{2}}\right)dt^{2}+\frac{dr^{2}}{\left(r^{2}+1-\frac{r_{0}^{2}}{r^{2}}\right)}+r^{2}\left(d\phi^{2}+\sin^{2}{\phi}\ d\Omega^{2}\right)+d\psi^{2}+sin^{2}\psi\ d\Omega^{2}_{4}.$ (273)
@@
$g_{\lambda\lambda}=g_{\lambda\theta}=0,\quad\text{and}\quad g_{\lambda\xi}=1.$ (274)
@@
$t(\lambda,\xi,\theta)=\int_{0}^{\lambda}\dot{t}\ d\lambda^{\prime}-\frac{\xi}{E}+J\theta\ ,\quad r(\lambda,\xi,\theta)=\int_{0}^{\lambda}\dot{r}\ d\lambda^{\prime}\ ,\quad\phi(\lambda,\xi,\theta)=E\theta+\int_{0}^{\lambda}\dot{\phi}d\lambda^{\prime}.$ (275)
@@
$+\left(-a_{0}J\dot{t}+JE\right)2d\lambda d\theta+(a_{0}JE^{-1})2d\xi d\theta+r^{2}\sin^{2}{\phi}\ d\Omega^{2}_{2}+ds^{2}(S^{5}),$
$=2d\lambda d\xi-a_{0}E^{-2}d\xi^{2}+\left(r^{2}E^{2}-J^{2}a_{0}\right)d\theta^{2}+Ja_{0}E^{-1}2d\xi d\theta+r^{2}\sin^{2}{\phi}.\ d\Omega^{2}_{2}+ds^{2}(S^{5})$
@@
$\lambda\rightarrow\lambda,\quad\xi\rightarrow\frac{\xi}{R^{2}},\quad\theta\rightarrow\frac{\theta}{R},\quad\psi\rightarrow\frac{\psi}{R},$ (276)
@@
$ds^{2}$ $=2d\lambda d\xi+\left(r^{2}E^{2}-J^{2}a_{0}\right)d\theta^{2}+r^{2}\sin^{2}\left(\int_{0}^{\lambda}\frac{J}{r^{2}}d\lambda^{\prime}\right)\ ds^{2}(\mathbb{R}^{2})+ds^{2}(\mathbb{R}^{5}),$ (277)
$=2d\lambda d\xi+{\sum\limits_{i=1}^{8}\left(C_{i}dy^{i}\right)^{2}}.$ (278)
@@
$ds^{2}$ $=2dx^{+}dx^{-}+\left[\sum\limits_{i}{\frac{C^{{}^{\prime\prime}}(x^{+})}{C_{i}}x^{i}x^{i}}\right]\left(dx^{+}\right)^{2}+\sum\limits_{i}dx^{i}dx^{i},$
@@
$\frac{d^{2}}{d\lambda^{2}}=\frac{d}{d\lambda}\left(\dot{r}\frac{d}{dr}\right)=\dot{r}^{2}\frac{d^{2}}{dr^{2}}+\dot{r}\left(\frac{d\dot{r}}{dr}\right)\frac{d}{dr}.$ (280)
@@
$M_{11}$ $=r^{-1}\frac{d}{dr}\left[\dot{r}\frac{d}{dr}\left(\dot{r}r\right)\right]=\frac{4J^{2}r_{0}^{2}}{r^{6}};\quad M_{44}=...=M_{88}=0,$
$M_{22}$ $=M_{33}=\frac{\dot{r}}{r\ \sin{\left(\int\frac{J}{r^{2}}du\right)}}\frac{d}{dr}\left[\dot{r}\frac{d}{dr}\left\{r\sin{\left(\int\frac{J}{r^{2}}du\right)}\right\}\right]=-\frac{2J^{2}r_{0}^{2}}{r^{6}}.$ (281)
@@
$ds^{2}=2dx^{+}dx^{-}+\left[\frac{4J^{2}r_{0}^{2}}{r^{6}}x_{1}^{2}-\frac{2J^{2}r_{0}^{2}}{r^{6}}(x_{2}^{2}+x_{3}^{2})\right]\left(dx^{+}\right)^{2}+\sum\limits_{i=1}^{3}dx_{i}^{2}+\sum\limits_{i=1}^{5}dy_{i}^{2}.$ (282)
@@
Now expanding the metric in the large R limit, we obtain the pp wave in the Rosen coordinates